OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (10, 12) and is perpendicular to 5x + 8y = 1 Please show all of your

OpenStudy (anonymous):

Step 1: put 5x + 8y = 1 into slope-intercept form (y=mx+b)

OpenStudy (anonymous):

8y = -5x + 1 y = -5/8x + 1/8 So your slope for this line is -5/8.

OpenStudy (anonymous):

12 = 8/5(10) + b 12 = 16 + b -4 = b So your equation is: y = 8/5x - 4

OpenStudy (anonymous):

Note that the slopes of the two lines are opposite reciprocals. I used the point given and the slope 8/5 to find the b value.